A four-stroke petrol engine has the following data:

      Number of cylinders=4

           Fuel used = 19.2 kg/hour

           Fuel to air ratio = 1:16 by weight

          Suction pressure = 1bar

          Suction temperature = 30oC

      Stroke to bore ratio = 1.25

      Engine speed = 2400rpm

       Volumetric Efficiency = 78%

Find the stroke length of engine[m]. Write the answer with 3 decimal places and do not add the unit.


1
Expert's answer
2021-11-30T14:16:01-0500

Mass flow rate of air (Ma) =19.2×16=307.2 kg/hr


Total rate of injection of air and fuel =307.2+19.2=326.4kg/hr


In kg/sec =. "\frac{326.4}{3600}=0.09067kg/s"


Mass flow rate "=\frac{Mass.N_e}{N_r}"


Where "N_e=" Engine speed

"N_r=" 2 for a four stroke engine


"N_e=\frac{2400}{60}=40rps"


Mass M "=\frac{0.09067×2}{40}=0.004534kg"



PV =MRT


Where P=1 bar ="1\u00d710^5Pa"


"R=" Gas constant =300J/kg/k

T= 30°c "=273.15+30=303.15"




"V=\frac{0.004534×300×303.15}{100000}\\=0.004123m^3"



Swept volume = "\frac{volume}{efficiency}"

Vd = "\frac{0.004123}{0.78}"


"=0.005286"


"V_d=\frac{ALN_{e}n}{n_r}"


Where A= cross section of cylinder

L= Stroke of cylinder

n= Number of cylinders



"AL=\frac{0.005286×2}{40×4}=6.608×10^{-5}"


Let D= bore

D =0.8L


"AL=\frac{\pi×0.64L^2.L}{4}=6.608×10^{-5}"



"L^3={1.315\u00d710^{-4}}"


"L=0.051"